Two 2.5 kg bowling balls are 25 m apart. What is the force of gravity exerted on each bowling ball? 


1
Expert's answer
2021-04-08T12:09:05-0400

"F_{12}=F_{21}=G\frac{m_1m_2}{r^2}="


"=6.67\cdot10^{-11}\cdot\frac{2.5×2.5}{25^2}=6.67×10^{-9}(N)=6.67(nN)"